OpenStudy (xxhomeworkhelperxx):

VERY EASY QUESTION FOR MEDALS FREE MEDALS Convert the Words into an Equation One More than One Equals Two

OpenStudy (xxhomeworkhelperxx):

Whoever Replys First

OpenStudy (anonymous):

1+1=2

OpenStudy (spyro517):

2=2

OpenStudy (anonymous):

if it is so easy why dont u do it -.-

OpenStudy (anonymous):

^

OpenStudy (xxhomeworkhelperxx):

Spyro your wrong I alreayd knwo the Answer

OpenStudy (xxhomeworkhelperxx):

I'm giving our Free Medals

OpenStudy (anonymous):

1 + 1 = 2

OpenStudy (anonymous):

Lol

OpenStudy (anonymous):

well damn i missed it

OpenStudy (anonymous):

already got it bud ;)

OpenStudy (spyro517):

I just put it differently. 1 more than 1 can also be 2 so I just put 2=2

OpenStudy (anonymous):

1+2-1=2

OpenStudy (anonymous):

spyro give me medal and i will giv u medal XD

OpenStudy (anonymous):

yay thanx

OpenStudy (spyro517):

medals for everyone :D